OpenStudy (anonymous):

Solve. x^2 – 4x = 0 a. x = 0 and x = –4 b. x = 2 and x = –2 c. x = 4 and x = 0 d. x = 0

myininaya (myininaya):

do you know how to factor x^2-4x

OpenStudy (anonymous):

yes, its x - 4, right?

myininaya (myininaya):

\[x(x-4)\] since \[x(x-4)=x \cdot x - x \cdot 4=x^2-4x\]

myininaya (myininaya):

so we have the following we need to solve \[x(x-4)=0\] which implies x=0 or x-4=0

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

and x = 4?

OpenStudy (anonymous):

is it?